### v3

# The Nontrivial Vacuum Structure of an Extended $t\overline{t}$ BEH (Higgs) Bound state

Abstract
In a recent reformulation of top-quark condensation for the Brout-Englert-Higgs boson, we introduced an extended internal wave-function, $\phi(r)$. We show how this leads to a manifestly Lorentz invariant formalism, where the absence of “relative time” is a gauge invariance of the bilocal field theory. This dictates a novel and nontrivial Lorentz invariant vacuum structure for the BEH boson, the relativistic generalization of a condensed matter state analogous to a BCS condensate.
preprint: Delafield-1220-2025
February 16, 2026
I Introduction
Beginning with Schrödinger, any two-body bound state can be described in a semiclassical limit as a bilocal field, $H(x,y)$ Schrodinger ; Yukawa . For the ground state ansatz this factorizes in barycentric (center-of-mass system) coordinates as:
$H(x^{\mu},y^{\mu})\sim H(X^{\mu})\phi(r^{\mu}).$ (1)
In the case of a Brout-Englert-Higgs (BEH) boson, composed of $\overline{t}t$, massless chiral top and anti-top quarks are located at space-time coordinates $(x^{\mu},y^{\mu})=X^{\mu}\pm r^{\mu}$. $H(X^{\mu})$ can be viewed as the standard model (SM) BEH isodoublet with electroweak charges and $\phi(r^{\mu})$ is a complex scalar which is electroweak neutral chris1 ; chris2 . $H(X^{\mu})$ describes the center-of mass motion of the BEH boson, and $\phi(r^{\mu})$ is then the internal wave-function.
A theory of a BEH boson composed of $\overline{t}t$, known as “top condensation,” was proposed in the 1990’s Yama ; BHL ; Topcolor ; NSD , deploying the Nambu-Jona-Lasinio (NJL) model NJL with renormalization group (RG) improvements BHL . The NJL model, however, is pointlike, lacking $\phi(r^{\mu})$, which leads to difficulties when there is a large hierarchy between the composite scale, $M_{0}$, and the electroweak scale $|\mu|$ (the symmetric phase mass of the BEH boson, $|\mu|=88$ GeV). The inclusion of the internal wave-function yields significant improvement in the predictions of the low energy parameters. In the limit $|\mu|<\!\!<M_{0}$ there is significant wave-function spreading of $\phi(\vec{r})$ and the resulting dilution effects dominate the low energy effective theory. This brings its predictions into concordance with experiment, virtually eliminating fine-tuning, and predicting the new mass scale of the binding interaction, $M_{0}\sim 6$ TeV, ref.(chris1 ,chris2 ) summarized in Appendix B.
We emphasize that the theory is manifestly Lorentz invariant. The challenge is, however, that the low energy internal wave-function, $\phi(r^{\mu})$, introduces a priori unwanted dependence upon “relative time.” This is $r^{0}=(x^{0}-y^{0})/2$ in the rest frame of the bound state, but boosted relative time would occur in any frame. Due to the single time parameter of Hamiltonian based quantum mechanics, the relative time is unphysical in a wave-function.
The wave-function can be viewed as the “end-cap” of the path integral that begins (and terminates) on given time-slices. The path-integral is a Green’s function of the Hamiltonian–Schrödinger equation that propagates the wave-function from an initial time and spatial configuration $(t_{0},\vec{X}_{0},\vec{r})$ to a future $(t_{1},\vec{X}_{1},\vec{r}_{1})$ (this is true in field theory where the initial and final configurations are static fields, $\phi(\vec{x}_{i})$, specified on given time slices, $t$). The relative time does not exist on these end-cap time slices and plays no role in the initial data. [^1] We remind the reader that, for the Hydrogen atom with potential, $V(\vec{r})$ and Schrödinger equation $H\psi(\vec{r},t)=i\partial_{t}\psi(\vec{r},t)$, the parameters $t$ and $\vec{r}$ do not form a 4-vector under Lorentz covariance. Rather, $(t,\vec{X})\rightarrow X^{\mu}$ becomes a relativistic 4-vector, with $\vec{X}$ describing the center-of-mass of the atom, and $\vec{r}\rightarrow r^{\mu}$ with the relative time $r^{0}$ constrained to zero. Hence, for wave-functions, relative time must be removed in a manner consistent with Lorentz invariance. The issue of relative time is avoided in the NJL model due to the pointlike interaction, but the absence of relative time is a well known challenge known to arise in any bound state with an extended interaction, Dirac .
The problem of maintaining Lorentz invariance is that relative time implies an “arrow of relative time,” a timelike, unit, 4-vector $\omega^{\mu}$, associated with $\phi(r^{\mu})$. The relative time is then $\tau$, where $r^{\mu}=\omega^{\mu}\tau$. The absence of relative time can then be viewed as a gauge symmetry of the internal wave-function $\phi(r^{\mu})$, where a gauge transformation is $r^{\mu}\rightarrow r^{\mu}+\omega^{\mu}\tau$ with $\tau$ acting as a gauge parameter. Given an $\omega^{\mu}$ we can then pass to a manifestly gauge invariant $\phi_{\omega}(r^{\mu})$ field, the analogue of a “Stueckelberg” field, where the symmetry is built in:
$\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})\equiv\phi\left(\omega^{\mu}\omega_{\nu}r^{\nu}-r^{\mu}\right)\qquad\omega^{2}\equiv\omega_{\mu}\omega^{\mu}=1.$ (2)
The issue then becomes how to avoid Lorentz non-invariance with an arbitrary input $\omega^{\mu}$, in particular, “what defines $\omega^{\mu}$?” For simple two-body bound states, where $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$, then $\omega^{\mu}$ can be identified with the normalized 4-momentum, $\omega^{\mu}=P^{\mu}/\sqrt{P^{2}}$ where $P^{2}=P_{\mu}P^{\mu}=\mu^{2}>0$. Hence, a two-body spherically symmetrical bound state becomes:
$H(X^{\mu})\phi_{\omega}(r^{\mu})=\exp(iP_{\mu}X^{\mu})\;\phi\!\left(\sqrt{((P_{\mu}r^{\mu})^{2}/P^{2})-r_{\mu}r^{\mu}}\right).$ (3)
This state is then explicitly Lorentz invariant. Then $\phi(r^{\mu})$ reverts to $\phi(|\vec{r}|)$ in the rest frame where it can be treated as a solution to a static Schrödinger-Klein-Gordon (SKG) equation chris1 .
However, there remains the question when $\mu^{2}<0$ and spontaneous symmetry breaking occurs: “How can we have a Lorentz invariant vacuum state with $P^{\mu}=0$, but nonzero $\omega^{\mu}$?” That is, “what determines $\omega^{\mu}$ in the vacuum?” Any constraint that locks $P^{\mu}$ to $\omega^{\mu}$ in the two-body case ceases to exist in the vacuum and $\omega^{\mu}$ would seem to become arbitrary. While one possibility is that $\phi(r^{\mu})\rightarrow$(constant) in the vacuum, this does not lead to a consistent solution to the SKG equation.
It would seem nonsensical to assert that the vacuum is defined by a condensate that occurred in a particular Lorentz frame with a random $\omega^{\mu}$. If, for example, $\omega^{\mu}$ is somehow associated with the local cosmic rest frame we would obtain induced Lorentz violating effects in the electroweak physics, e.g., vacuum Cerenkov radiation for all particles that receive mass from the BEH boson, top-quarks to electrons and neutrinos ColemanGlashow . We have estimated these effects in chris1 and they are suppressed as $\sim|\mu|^{2}/M_{0}^{2}$. While the limit on vacuum Cerenkov radiation for the electron is satisfied, these effects may be problematic, e.g., potentially large radiatively induced Lorentz non-invariant corrections to electrodynamics may arise at loop level Kostelecky . We therefore require a starting point in which the vacuum is manifestly Lorentz invariant, hence it must contain no preferred $\omega^{\mu}$. Yet, we evidently require nonzero $\omega^{\mu}$ to define the solutions, $\phi_{\omega}(r^{\mu})$, to implement the relative time invariance.
Hence, we propose the following solution to the vacuum problem: the vacuum is a Lorentz invariant sum over all frames of the individual solutions $\phi_{\omega}(r^{\mu})$ in each frame. We introduce a Lorentz invariant integral over $\omega^{\mu}$, leading to a novel internal wave-function for the vacuum state, $\Phi(r^{\mu})$, with (unrenormalized) $H^{\prime}(X)$.
$H(x,y)=H^{\prime}(X)\Phi(r^{\mu})\qquad\text{where,}\qquad\Phi(r^{\mu})={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu}).$ (4)
With this definition of the vacuum, $\Phi(r^{\mu})$ becomes a collective state, similar to a condensed matter system such as the BCS superconductor.
Our present calculation is formally similar to the construction of the coherent condensate of Cooper pairs in a BCS superconductor Cooper BCS , differing essentially by our required Lorentz invariance. Cooper pairs are two-body bound states, (approximate) bilocal plane waves in momentum space: $\phi_{i}(\vec{r})\approx\phi_{i}\exp(\vec{k}_{i}\cdot(\vec{x}-\vec{y}))$, where the $\vec{k}_{i}$ lie in a small range immediately above the Fermi surface of the material, hence the “pairing” is an electron of momentum $\vec{k}_{i}$ with the antipodal electron $\vec{k}_{i}$. The Hamiltonian kinetic term for the pairs is then $\small\sim(1/m)\sum_{i}|\phi_{k_{i}}|^{2}k^{2}_{i}\propto N$, while the two-body attractive scattering potential (weak phonon exchange interaction) is $\small\sim-\epsilon m\sum_{i}\sum_{j}\phi^{\dagger}_{k_{i}}\phi_{k_{j}}\propto N^{2}$. Hence the weak phonon interaction is significantly enhanced by large $N\epsilon\sim 1$ and can compete with the kinetic terms to form a stable collective state which forms the condensate. The semiclassical wave-function of the ground state is then $\Phi\sim\sum_{i}^{N}\phi_{i}(\vec{r})$ where the sum spans the Fermi surface. The BCS theory is essentially a bilocal field theory. [^2] The BCS ground state is a coherent state of Cooper pairs, and specified as a product of Dirac kets that are mixtures of vacuum and pairs, $|BCS\rangle\sim\prod^{N_{F}}_{i}(u_{i}|0\rangle+v_{i}|\!\uparrow\!\vec{k}_{i},\!\downarrow\!-\vec{k}_{i}\rangle)$, where the product extends over the $N_{F}$ momenta, $k_{i}$, slightly above the Fermi surface (bounded above by the Debye frequency $\omega_{D}$). The superconductor’s properties are determined by the resulting ratio $v_{i}/u_{i}$, obtained by minimization of the Hamiltonian with, e.g., temperature effects or current dependence, etc. (See e.g., Feynstat Urbana , and related discussion of mean-field approximations GL BEC ). The condensate quantum field is bilocal, with creation operators $a^{\dagger}_{\vec{k}_{i}}$ for electrons of momenta $\vec{k}_{i}$, $\widehat{\Phi}(\vec{r})\sim\sum^{N}_{i}a^{\dagger}_{\vec{k}_{i}}a^{\dagger}_{-{\vec{k}_{i}}}\phi_{i}(\vec{r})$. Note this creates, and sums over, Cooper pairs of antipodal momenta $(\vec{k}_{i},-\vec{k}_{i})$ that span the Fermi surface. The analogue of our semiclassical bilocal field is then $\Phi(\vec{r})=\langle BCS|\widehat{\Phi}(\vec{r})|BCS\rangle\sim\sum^{N}_{i}u_{i}v_{i}^{*}\phi_{i}(\vec{r})$. The ground state (vacuum) expectation value of $\langle\Phi\rangle$ is determined by a gap equation (similar of the original formulation of the NJL model NJL ). A gap equation is equivalent to the minimization of the BEH (Higgs) potential (the “quartic interaction” of the potential is implicit in the gap equation loop, and for us the quartic coupling also arises at loop level).
The Brout-Englert-Higgs (BEH) vacuum wave function which we presently propose, $\Phi(r^{\mu})$, is a sum over all constituent wave-function solutions of the SKG equation, each with $P^{\mu}=0$ but with nontrivial $\phi_{\omega}(r^{\mu})$, each corresponding to a different Lorentz frame, $\omega^{\mu}$, the analogue of the Fermi momenta in the superconductor. The $\phi_{\omega}(r^{\mu})$ are then a relativistic generalization of Cooper pairs BCS , but while Cooper pairs span the Fermi surface in a superconducting condensate, the $\Phi\sim\int_{\omega}\phi_{\omega}$ span the future timelike hyperboloid in the timelike unit 4-vectors, $\omega^{\mu}$.
The component fields, $\phi_{\omega}(r)$, each satisfy a nontrivial integro-differential equation with dependence upon the frame $\omega^{\mu}$. It follows that the invariant integral over $\omega^{\mu}$, normalized by $\cal{N}$}, yields a Lorentz invariant $\Phi(r^{\mu})$. The Hamiltonian is then diagonalized by the coherent state $H(X)\Phi(r^{\mu})$. The normalized effective action for $H(X^{\mu})$ is obtained and, upon integrating out $r^{\mu}$, yielding the “Higgs” potential $\mu^{2}|H|^{2}+(\lambda/2)|H|^{4}$. $H$ then acquires a vacuum expectation value (VEV), determined by a negative SKG eigenvalue, $\mu^{2}$, and stabilized by the loop induced quartic interaction $\lambda$.
Figure 1: Vacuum wave-function, $\Phi(r^{\mu})$, spanning the timelike hyperboloid in 4-vectors $\omega^{\mu}$ by integrating over “internal wave-functions,” $\phi_{\omega}(r)$, to form the Lorentz invariant $\Phi(r^{\mu})={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})$. This is a relativistic analogue of a BCS state that integrates over the Cooper pairs on the Fermi surface.
The vacuum emerges from the underlying theory with $H^{\prime}(X)\Phi(r^{\mu})=H(X^{\mu})\widetilde{\Phi}(r)$ where $\widetilde{\Phi}(r)=(1/\sqrt{N})\Phi(r^{\mu})$ is the classical average over all $\phi_{\omega}(r^{\mu})$ and $\langle H\rangle=v_{weak}$. The BEH boson observed at the LHC, $h(X)$ and the Nambu–Goldstone phases that become longitudinal $W^{\pm}$ and $Z^{0}$, then emerge as “excitons” of the collective state. In the broken phase we then have:
$H(X^{\mu})\widetilde{\Phi}(r^{\mu})\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}+\frac{1}{\sqrt{2}}h(X)\\ 0\end{array}\biggr)\widetilde{\Phi}(r^{\mu}),$ (7)
where $h(X)$ is the “Higgs field” of the standard model.
Here the constant zero 4-momentum VEV, $v_{weak}$, is carried by $H(X^{\mu})$ and determined in the usual way by the minimum of the sombrero potential. The main prediction of the theory remains as the existence of the new binding interaction at the scale $M_{0}\approx 6$ TeV and the emergence of an octet of colorons coupled most strongly to third generation quarks. This interaction is only partially strong (approximately half critical) since the critical behavior occurs only in the binding channel where loop effects reinforce the binding interaction chris2 Jackiw .
In the present paper we will mainly focus on the nontrivial vacuum. We will also briefly sketch how higher dimension operator terms ${\cal{O}}(1/M_{0}^{2})$ may be extracted from the Yukawa coupling in Section IV (see also chris1 ), which may ultimately present observables in high sensitivity flavor physics experiments. We mainly ignore the complications of the introduction of all flavors of quarks and leptons, in particular the $b_{R}$ quark. We expect these to be perturbative and follow the earlier papers on extended technicolor Technicolor and “topcolor” from the 1990’s Topcolor , and will be the subject of future work christodo .
Our main result is a successful, natural, minimally fine-tuned (few $\%$), composite theory of the BEH boson, with a novel physical scale $\sim 6$ TeV corresponding to a new semi-strong interaction within the third generation quarks, and associated gauge fields that are possibly accessible to the LHC (chris1 ,chris2 ). We show that vacuum of this theory maintains the Lorentz invariance, as a dynamical analogue of a BCS superconductor, yielding the spontaneous symmetry breaking seen in the SM and a composite BEH boson.
II Implementing the Relative Time Symmetry
We begin with the notion that the internal wave-function of two-body bound state, $\phi(r^{\mu})$, must not depend upon “relative time,” i.e., is independent of $r^{0}$ in the barycentric frame, or correspondingly $r^{\prime 0}$ in any frame. This can be seen in the free field limit by kinematics if we consider a pair of equal-mass particles of 4-momenta $p_{1}$ and $p_{2}$, $p_{1}^{2}=p_{2}^{2}=m^{2}$, and a bilocal wave-function, $H(x,y)\sim\exp(ip_{1}x+ip_{2}y)$. We pass to the total momentum $P=(p_{1}+p_{2})$ and relative momentum $Q=(p_{1}-p_{2})$, and the plane waves become $\exp(iPX+iQr)$, where we define “barycentric coordinates,” $x^{\mu}=X^{\mu}+r^{\mu}$, $y^{\mu}=X^{\mu}-r^{\mu}$. Note that $P_{\mu}Q^{\mu}=p_{1}^{2}-p_{2}^{2}=0$. Therefore, in the center-of-mass frame, in which $P=(P^{0},\vec{0})$ and $Q=(0,\vec{q})$, we see that $Q^{0}=0$. This implies there is no dependence in the bilocal state on $r^{0}$ through $\exp(iQ_{0}r^{0})=1$, and likewise no dependence upon a boosted relative time $r^{\prime 0}$ in any other frame. If the particles are constituents of a bound state, then the “relative time” must decouple from the dynamics. Indeed, it is useful to think of the bound state as free field pair of particles for which the bilocal field is properly normalized, and the interaction is subsequently adiabatically switched on.
Given an arbitrary Lorentz invariant function, $\phi(r^{\mu})$, in any frame there will generally be dependence upon a relative time, $\sim\tau$. This is analogous to the gauge dependent components of a vector potential, and it is an artifact of using the bilocal field description. [^3] I am grateful to Bill Bardeen for some discussions that inspired this perspective. As described in the Introduction, the relative time, $\tau$, can be written in any given frame as $r^{\mu}=\omega^{\mu}\tau$, and implicitly requires the timelike, 4-vector, $\omega^{\mu}$, the “arrow of relative time.” Hence, eliminating dependence upon $\tau$, requires a Lorentz invariant constraint, such as $\omega_{\mu}\partial^{\mu}\phi(r)=0$. In the symmetric phase of the standard model (SM), (or for any typical two-body bound state) the BEH boson contains such a vector, i.e., the 4-momentum $P^{\mu}$ carried by $H(X^{\mu})\sim\exp(iP_{\mu}X^{\mu})$. We can therefore bootstrap $\omega^{\mu}$ to $P^{\mu}$ through a constraint relation $\omega^{\mu}\propto P^{\mu}$. For example, we can do this semiclassically by introducing Lagrange multipliers into the action, such as:
$W=M_{0}^{4}\int d^{4}Xd^{4}r\;\lambda^{\prime}(H^{\dagger}D_{\mu}H\phi^{\dagger}\partial_{\mu}\phi),$ (8)
where we demand $\delta W/\delta\lambda^{\prime}=0$, which imposes the kinematic constraint $P_{\mu}Q^{\mu}=0$. However, the question then remains: “what happens in the vacuum where $P_{\mu}=0$ and $\omega^{\mu}$ becomes unconstrained?”
Consider the bilocal field in barycentric coordinates:
$H(x,y)\rightarrow H(X^{\mu})\phi_{\omega}(r^{\mu}).$ (9)
We will presently focus upon the fields, $\phi_{\omega}(r^{\mu})$, as defined in eq.(2) which are invariant under the gauge transformation $r^{\mu}\rightarrow\omega^{\mu}\tau$ and thus have no dependence upon $\tau$, though an implicit dependence upon $\omega^{\mu}$ remains. Indeed, $\phi_{\omega}(r^{\mu})$ is then the analogue of a “Stueckelberg” field, e.g., a gauge field such as $B_{\mu}=A_{\mu}-\partial_{\mu}\chi$ which is invariant under $A_{\mu}\rightarrow A_{\mu}+\partial_{\mu}\tau$ and $\chi\rightarrow\chi+\tau$. Note that eq.(2) satisfies the constraint equation $0=\omega^{\mu}\partial_{\mu}\phi(r^{\mu})$. In the following, $\phi(r^{\mu})$ will refer to an arbitrary Lorentz invariant scalar, while $\phi_{\omega}(r^{\mu})$ is of the (Stueckelberg) form of eq.2) with the relative time projected out.
Consider the action $S_{\phi}$ for the internal field $\phi(r^{\mu})$ that arises in natural top condensation chris1 , of eq.(87). We replace $\phi(r^{\mu})\rightarrow\phi_{\omega}(r^{\mu})$ in $S_{\phi}$,
$\!\!\!\!\!\!\!\!\!\!S_{\phi}$ $\rightarrow$ $M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr),$ (10)
where $\partial_{\mu}=\partial/\partial r^{\mu}$ and $D_{F}(2r)$ is defined in eq.(89).
If we then vary $S_{\phi}$ with respect to $\phi_{\omega}+\delta\phi_{\omega}$ we obtain a formal, manifestly Lorentz invariant integro-differential equation:
$M_{0}\int\!\!\biggl(-Z\frac{\partial^{2}\phi_{\omega}(r^{\mu})}{\partial r^{\mu}\partial r_{\mu}}+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega}(r^{\mu})\biggr)\omega_{\nu}dr^{\nu}=ZM_{0}\int\!\!\mu^{2}\phi_{\omega}(r^{\mu})\omega_{\nu}dr^{\nu}=\mu^{2}\phi_{\omega}(r^{\mu}).$ (11)
Note the presence of the overall line integral, $\int\omega_{\mu}dr^{\mu}$. This line integral remains in the equation of motion since $\phi_{\omega}(r^{\mu})$ has no dependence upon $r_{\mu}\propto\omega_{\mu}$, hence the variation is constrained, $\delta\phi_{\omega}\sim\delta^{3}(r^{\mu}_{\perp})$ where $\omega_{\mu}r^{\mu}_{\perp}=0$ and does not produce a longitudinal variation, $\delta(\omega_{\mu}r^{\mu}$).
We define $Z$ by the line integral normalization chris1 ,
$ZM_{0}\int dr^{\mu}\omega_{\mu}=1.$ (12)
$Z$ is defined to canonically normalize the free field pair of particles before the interaction is turned on. It is important to realize that eq.(11) is not a conventional Klein-Gordon equation due to the line integral constraint.
We can find solutions to eq.(11) as follows: Since $S_{\phi}$ is Lorentz invariant we can evaluate the action in the particular frame, $\omega^{\mu}=(1,0,0,0)$, where $\phi_{\omega}(r^{\mu})\rightarrow\phi(0,\vec{r})\equiv\phi(\vec{r})$ and hence $ZM_{0}\int dr^{\mu}\omega_{\mu}\rightarrow ZM_{0}\int dr^{0}=1$. In this frame the action becomes:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{3}\!\!\int\!\!d^{3}r\;\biggl(-|\nabla_{\vec{r}}\phi(\vec{r})|^{2}+\int dr^{0}\;{2g_{0}^{2}N_{c}M_{0}}D_{F}(2r^{\mu})|\phi(\vec{r})\;|^{2}\biggr).$ (13)
Using eq.(12) converts the normalization of eq.(82) to:
$M_{0}^{4}Z\int\!\!d^{4}r|\phi(r^{\mu})|^{2}\rightarrow\int\!\!d^{3}r\;M_{0}^{3}\;|\phi(\vec{r})|^{2}=1,$ (14)
and yields, in this frame, the Yukawa potential chris1 :
$V(2|\vec{r}|)=\int dr^{0}2g_{0}^{2}N_{c}D_{F}(2r^{\mu})=-\frac{g_{0}^{2}N_{c}e^{-2M_{0}|\vec{r}|}}{8\pi|\vec{r}|}.$ (15)
The SKG equation in the spherical ground state thus becomes:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!-\nabla^{2}\phi(r)-g_{0}^{2}N_{c}M\frac{e^{-2M_{0}r}}{8\pi r}\phi(r)=\mu^{2}\phi(r)\qquad\nabla^{2}=\biggl(\frac{\partial^{2}}{\partial r^{2}}+\frac{2}{r}\frac{\partial}{\partial r}\biggr)\qquad r\equiv\sqrt{\vec{r}^{2}}.$ (16)
We study its properties and solutions in ref.chris1 .
In the solution to the SKG equations with the eigenvalue, $\mu^{2}$, we see upon integrating by parts that $S_{\phi}\rightarrow\mu^{2}$, and the full bound state action then becomes (in the pointlike limit of the interaction):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X^{\mu})|^{2}-\mu^{2}|H(X^{\mu})|^{2}-\frac{\lambda}{2}(H^{\dagger}H)^{2}-g_{Y}\left([\overline{\psi}_{iL}(X)t_{R}(X)]_{f}H^{i}(X)+h.c.\right)\biggr)+S^{\prime}.$ (17)
$\mu^{2}$ is the physical mass of the bound state. The Yukawa coupling, $g_{Y}$ and quartic coupling $\lambda$ are defined in eqs.(85-90) in Appendix B. We emphasize that, while we have evaluated the action in the particular $\omega^{\mu}$ frame in eq.(13), this is just a calculation in a simplifying frame in the overall Lorentz invariant action eq.(11). The result eq.(17) holds in any frame.
As shown in ref.chris1 , the SKG equation has a critical coupling, $g_{0}=g_{c}$, for which $\mu^{2}=0$, very close to the quantum NJL critical coupling:
$\frac{g_{c}^{2}N_{c}}{8\pi^{2}}=1.06940,\qquad\makebox{c.f, the NJL critical value, }\qquad\left.\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right|_{NJL}=1.00.$ (18)
The loop level (NJL-like) effects generate $\lambda$ and also add to the formation of the bound state as discussed in chris2 . This amplifies the coupling strength in the bound state channel and generates a renormalized coupling for the 4-fermion interaction, $\overline{g}^{2}_{0}$, where:
$\overline{g}_{0}^{2}=g_{0}^{2}\left(1-\frac{g_{0}^{2}N_{c}}{8\pi^{2}}\right)^{-1}.$ (19)
When $\overline{g}_{0}>g_{c}$ the eigenvalue is $\mu^{2}<0$. In such a solution the action eq.(17) for $H(X^{\mu})$ then yields the “sombrero potential”:
$\mu^{2}|H|^{2}+\frac{\lambda}{2}(H^{\dagger}H)^{2}\qquad\text{where,}\;\;\;\mu^{2}<0.$ (20)
The solution of the SKG equation for $\phi(r)$ can be obtained approximately analytically, or by numerical integration chris1 ; chris2 . At short distances $\phi(r)\sim\phi(0)$, and extends at large distances in the rest frame to, $\phi(r)\sim ce^{-|\mu|r}/r$, where $|\mu|<\!\!M_{0}$ and we are near critical coupling. The solution is normalized as in eq.(14), which dilutes the value of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$ and suppresses the Yukawa coupling $g_{Y}\propto\phi(0)$ and $\lambda\propto g_{Y}^{4}\propto|\phi(0)|^{4}$.
III Broken Phase, Manifestly Lorentz Invariant Vacuum, and BEH Excitations
Presently we show how the action can be written in terms of the collective field $\Phi(r^{\mu})$. We find it conceptually useful to begin by approximating an integral representation of the collective state by a discrete sum over $N$ of the solutions, $\omega_{i\mu}$. This approximates the continuous integrals which are defined subsequently corresponding to the large-$N$ limit. We begin with a brief discussion of the orthogonality of the $\phi_{\omega_{i}}$ which is further treated in Appendix A. In Appendix B for reference, we give a formal summary of the natural top condensation scheme.
III.1 Formal Derivation
The action of eq.(10) is manifestly Lorentz invariant with dependence upon the 4-vector $\omega^{\mu}$. For the spherically symmetric ground state it must yield a Lorentz invariant expression in $\omega$. Hence, under a Lorentz transformation, $\omega^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}\omega^{\nu}$ we can likewise perform $r^{\prime}_{\mu}=\Lambda_{\mu}^{\nu}r^{\nu}$, however, we are then free to change the integration variable back to its original form, $r^{\prime}\rightarrow r$ to obtain:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}=M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi_{\omega^{\prime}}^{\dagger}(r)\partial^{\mu}\phi_{\omega^{\prime}}(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi_{\omega^{\prime}}^{\dagger}(r)\phi_{\omega^{\prime}}(r)\biggr).$ (21)
We can see that there is orthogonality of the $\phi_{\omega}(r^{\mu})$ solutions in the kinetic term (or in integrating over any extended smooth, approximately constant, Lorentz invariant function of $r^{\mu}$). The orthogonality breaks down, however, in the pointlike limit of the interaction where the $\phi_{\omega}(r)$ will freely mix with any $\phi_{\omega^{\prime}}(r)$ where $\omega^{\prime}\neq\omega$.
Formally, for a smooth Lorentz invariant function, $F(r^{\mu})$ (or differential operator, e.g, $F\sim\partial^{2}$):
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(\omega-\omega^{\prime})\phi_{\omega}^{\dagger}(r^{\mu})F(r^{\mu})\phi_{\omega`}(r^{\mu}).$ (22)
If however, $F\sim\delta^{4}(r^{\mu})$ we have mixing of $\omega$ and $\omega^{\prime}$:
$ZM_{0}^{4}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r^{\mu})\delta^{4}(r^{\mu})\phi_{\omega`}(r^{\mu})=ZM_{0}^{4}\phi_{\omega}^{\dagger}(0)\phi_{\omega`}(0)=ZM_{0}^{4}\int\!\!d^{4}r\;\delta^{4}(r^{\mu})\phi_{\omega}^{\dagger}(r)\phi_{\omega`}(r).$ (23)
This implies that we will have the BCS or BEC-like behavior for $\Phi(r^{\mu})$. We derive the orthogonality in Appendix A.
Define a collective field by summing over a large set of $N$ arbitrary $\omega_{i\mu}$ unit 4-vectors that span the future timelike hyperboloid:
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu}),\qquad\text{where,}\qquad\phi_{\omega_{i}}(r^{\mu})=\phi(\omega_{i}^{\mu}\omega_{i}^{\nu}r_{\nu}-r^{\mu}),\qquad\omega_{i}^{2}=1.$ (24)
The $\phi_{\omega_{i}}(r)$ each have an internal $\omega_{i\mu}$ and each is a solution of the equation of motion, eq.(11). We will use $\phi_{\omega}$ to represent any arbitrary $\omega$ field. In this framework the classical average of the $\phi_{\omega}$ is:
$\widetilde{\Phi}(r^{\mu})=\frac{1}{\sqrt{N}}\Phi(r^{\mu})=\frac{1}{{N}}\sum^{N}_{i=1}\phi_{\omega_{i}}(r^{\mu});\qquad\widetilde{\Phi}(r^{\mu})=\phi_{\omega_{i}}(0).$ (25)
Consider the action for the single two-body bound state $H(X^{\mu})\phi(r)$ of eqs.(81 to 85), replacing $\phi(r)\rightarrow\Phi(r^{\mu})$ and $H(x)\rightarrow H^{\prime}(X)$. and denote renormalized parameters by primes . For the sake of discussion we break the action into separate components:
$S=S_{1}+S_{2}+S_{3}+S_{Y}+S_{\lambda},\qquad\text{where,}\qquad H(X^{\mu})\phi(r)\rightarrow H^{\prime}(X)\Phi(r^{\mu}),$ (26)
where:
$S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r^{\mu})|^{2}\biggr),$ (27)
$S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr),$ (28)
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu})|H^{\prime\dagger}H^{\prime}||\Phi(r^{\mu}))|^{2}\biggr),$ (29)
and we have taken the pointlike limit of $D_{F}(2r^{\mu})$ of eq.(89),
$D_{F}(2r^{\mu})\rightarrow\frac{1}{M_{0}^{2}}\delta^{4}(2r^{\mu})=\frac{1}{16M_{0}^{2}}\delta^{4}(r^{\mu}).$ (30)
The Yukawa interaction is,
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}^{\prime}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.},\qquad\widehat{g}^{\prime}_{Y}\approx g^{\prime}{}_{0}^{2}\sqrt{2N_{c}/J},$ (31)
and the quartic interaction is given in the point-like limit of $\Phi(r^{\mu})$:
$\!\!\!\!\!\!\!\!\!\!S_{\lambda}=-\frac{\widehat{\lambda}}{2}\int\!\!d^{4}X(H^{\dagger}H)^{2}|\Phi(0)|^{4}{+h.c.}.$ (32)
Our problem is to verify that $S(H^{\prime}\Phi)$, with renormalized parameters, is consistent with the underlying theory of the $\phi_{\omega}=\phi(r)$ as defined in eqs.(81,87). We therefore substitute the collective field definition of eq.24 and define the renormalized parameters.
The orthogonality of the $\phi_{\omega_{i}}$ fields implies in a discrete sum:
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r),$
$\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega_{i}}^{\dagger}(r)\partial^{\mu}\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r).$ (33)
We use these relations and the following normalizations:
[…]
Note that these normalizations differ from the BCS superconductor in two major regards. First, $Z$ is treated as an extensive parameter, i.e., when we renormalize each $\phi_{\omega}\rightarrow(1/\sqrt{N})\phi_{\omega}$ to comprise the sum over $\omega^{\mu}$ we then redefine $Z\rightarrow Z^{\prime}=ZN$. The new composite bilocal field then takes the form $H^{\prime}(X)\Phi(r^{\mu})$.
Secondly, the $H^{\prime}(X)$ kinetic term should be canonical with the Lorentz invariant $\Phi(r^{\mu})$, hence we renormalize the $H^{\prime}(X)$ field as:
$H^{\prime}=H/\sqrt{N}\qquad\text{and then}\qquad Z^{\prime}={N}Z.$ (38)
This choice preserves the product $Z^{\prime}H^{\prime\dagger}H^{\prime}=ZH^{\dagger}H$, and $Z$ is still defined by the line integral relation of eq.(12) for the underlying $\phi_{\omega}$ field. These normalizations will yield the same predictions for $g^{2}_{critical}$ and $M_{0}\sim 6$ TeV, as one would obtain for a single $\phi_{\omega}$ condensate.
We now discuss the calculation of the full action term by term:
$S_{1}$: We have from eqs.(24,27):
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|DH^{\prime}(X)|^{2}|\Phi(r^{\mu})|^{2}\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\sum^{N}_{j=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH^{\prime}|^{2}\biggl(\frac{Z^{\prime}}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH|^{2}\biggl(\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;\text{for any }\omega,$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|DH|^{2}|\widetilde{\Phi}|^{2}=\!\!\int\!\!d^{4}X|DH|^{2},$ (39)
where we have:
$ZM_{0}^{4}\int\!\!d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2}=ZM_{0}^{4}\int\!\!d^{4}r|\phi_{\omega}(r^{\mu})|^{2}\;(\text{for any }\omega)=1.$ (40)
$S_{1}$ is therefore consistent with the underlying two-body theory. Note that $\widetilde{\Phi}_{\omega}$ and $\phi_{\omega}$ integrate out with the normalization eq.(40).
$S_{2}$: Likewise the $\Phi$ kinetic term becomes:
$\!\!\!\!\!\!\!\!\!\!S_{2}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z^{\prime}|H^{\prime}(X)|^{2}|\partial_{r}\Phi(r^{\mu})|^{2}\biggr)=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\phi_{\omega}^{\dagger}(r)\partial^{\mu}\phi_{\omega}(r)\biggr)\;(\text{for any }\omega)$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(|H|^{2}\partial_{\mu}\widetilde{\Phi}^{\dagger}(r)\partial^{\mu}\widetilde{\Phi}(r)\biggr),$ (41)
yielding consistency with the underlying kinetic term.
$S_{3}$: The interaction becomes:
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(2{g^{\prime}{}_{0}^{2}N_{c}}D_{F}(2r^{\mu})|H^{\prime}|^{2}|\Phi(r^{\mu}))|^{2}\biggr)\approx M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;{|H^{\prime}|^{2}}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{N}\sum^{N}_{i=1}\phi_{\omega_{i}}^{\dagger}(r)\sum^{N}_{j=1}\phi_{\omega_{j}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;N|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)\biggr)\;(\text{for any }\omega).$ (42)
The essential result is that the potential is approximately $\sim\delta^{4}(r^{\mu})$, hence the line integral orientation $\omega^{\mu}$ becomes irrelevant in the potential. In the second-to-last term we see the usual enhancement factor, $N$, that would normally lead to the BEC or BCS phenomena. However, in the last term we see that the normalization of $H^{\prime}=H/\sqrt{N}$ undoes the $N$-fold enhancement and the coupling is not renormalized $g_{0}^{\prime 2}=g_{0}^{2}$. This is therefore different than the case of the BCS superconductor or the BEC where the coupling is enhanced by a factor of $N$. The nonrenormalization of $g_{0}^{2}$ owes to the bilocal field theory with the renormalization of $H^{\prime}=H/\sqrt{N}$.
We can rewrite the interaction in the unprimed parameters as:
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r|H(X^{\mu})|^{2}\biggl(\frac{2g_{0}^{2}N_{c}}{M_{0}^{2}}D_{F}(2r)|\widetilde{\Phi}(r)|^{2}\biggr).$ (43)
To a good approximation we can freely swap between $D_{F}(2r)\leftrightarrow\delta^{4}(r)/16M_{0}^{2}$.
$S_{Y}$: The Yukawa interaction is:
$\!\!\!\!\!\!\!\!\!\!S_{Y}=g^{2}_{0}\sqrt{2JN_{c}}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{\prime i}(X)\Phi(\vec{r}){+h.c.}.$
$\approx g^{2}_{0}\sqrt{2N_{c}/J}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}\;\delta^{4}(r)\;H^{i}(X)\widetilde{\Phi}(r^{\mu}){+h.c.}.$
$=g_{Y}M_{0}^{2}\int\!\!d^{4}X\;[\overline{\psi}_{iL}(X)\psi_{R}(X)]_{f}\;\;H^{i}(X)\widetilde{\Phi}(0){+h.c.};\qquad\qquad g_{Y}=g_{0}^{2}\sqrt{2N_{c}/J},$ (44)
where we have taken the limit $D_{F}(2r)\rightarrow\delta^{4}(r)/16M_{0}^{2}$.
We see that $g_{Y}$ is nonrenormalized. The nonrenormalization of $g_{Y}$ implies that the result obtained for $M_{0}$ from the solution to the SKG equation remains intact, i.e., $M_{0}\sim 6$ TeV with $|\mu|\sim 88$ GeV. If we don’t take the strict $\delta$-function limit we can do an expansion of the integrand in $r^{\mu}$. This generates a series of higher dimension operators in inverse powers of $M_{0}^{2}$ providing potentially sensitive probes to $M_{0}$ and the shape of the wave-function $\Phi(r^{\mu})$, briefly described in section IV.
$S_{\lambda}$: The quartic term likewise involves,
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{1}{2}\hat{\lambda}^{\prime}\int d^{4}Xd^{4}r\;\delta^{4}(r)|H^{\prime}(X)\Phi(r^{\mu})|^{4}=\frac{1}{2}\hat{\lambda}\int d^{4}X|H(X^{\mu})|^{4}|\widetilde{\Phi}(0)|^{4},$ (45)
where $\lambda^{\prime}=\lambda$ is therefore not renormalized. Essentially $\lambda$ is determined by the Yukawa coupling, $g_{Y}$ and the renormalization group running $\lambda$ from $M_{0}$ to $|\mu|$ and the result we obtained from the underlying theory at one loop, e.g., $\lambda\approx 0.23$, is not modified. This term would also permit expansion in $r^{\mu}$ and yield an operator expansion of new physics.
Finally, the result of inserting solutions to the SKG equation, integrating the kinetic term by parts, yields the full action:
$S_{\Phi}=Z^{\prime}M_{0}^{4}\mu^{2}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\;|\widetilde{\Phi}|^{2}=\mu^{2}\!\!\int d^{4}X|H|^{2}.$ (46)
We thus obtain the same eigenvalue for $\Phi(r^{\mu})$ as for the underlying $\phi_{\omega}(r)$.
It is straightforward to repeat the above with a continuous, manifestly Lorentz invariant, integral representation. To match definitions used above we have to reintroduce $N$, and require the matching condition to the classical sum:
$\sum_{i}^{N}(c)={\cal{N^{\prime}}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\int_{\omega}(c)=N(c),$ (47)
where $c$ is an arbitrary constant. We then define the collective field, $\Phi(r^{\mu})$, and match the discrete and continuous representations,
$\Phi(r^{\mu})=\frac{1}{\sqrt{N}}\sum_{i}^{N}\phi_{\omega_{i}}={\cal{N}}\int d^{4}\omega\;\delta(\omega^{2}-1)\;\phi_{\omega}(r^{\mu})\equiv\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}(r^{\mu}).$ (48)
Hence ${\cal{N^{\prime}}}=\sqrt{N}{\cal{N}}$.
Note that $\cal{N}$ is the normalization of a divergent integral over the hyperboloid, and requires, in principle, regularization. We will not enter into a detailed discussion of regularization here, however we note that we can usually perform a “Wick rotation.” If the integral is an analytic function of the metric, $g_{\mu\nu}=(1,-1,-1,-1)$, it can be continued as $g_{\mu\nu}\rightarrow\eta_{\mu\nu}\sim(1,1,1,1)$. Then the hyperboloid is replaced by a Euclidean 4-sphere and
$N={\cal{N}}\int\!\!d^{4}\hat{\omega}\;\delta(1-\hat{\omega}^{2})={\cal{N}}\pi^{2}\int\hat{\omega}^{2}d\hat{\omega}^{2}\delta(\hat{\omega}^{2}-1)=\pi^{2}{\cal{N}},$ (49)
We then replace $\eta$ by $g$. Moreover, alternative averaging functions could be defined. For example, if we define $\phi_{\omega}(r^{\mu})=\phi(\omega^{\mu}(\omega\cdot r)-r^{\mu})$ then we could take $\int_{\omega}\rightarrow\int d^{4-\epsilon}\omega$ as in a momentum integral, and thus use dimensional regularization.
To apply this, consider the expression for the normalization of $H^{\prime}\Phi$, similar to the $S_{1}$ calculation above:
$Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}|\Phi(r^{\mu})|^{2}=Z^{\prime}M_{0}^{4}\int d^{4}Xd^{4}r|H^{\prime}|^{2}\frac{1}{\sqrt{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi^{\dagger}_{\omega^{\prime}}(r)$
$=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2}\frac{1}{{N}}\int_{\omega}\phi^{\dagger}_{\omega}(r)\phi^{\dagger}_{\omega}(r)\times\biggl(\int_{\omega^{\prime}}\biggr)=ZM_{0}^{4}\int d^{4}Xd^{4}r|H|^{2},$ (50)
where we use the orthogonality relation,
$\int\!\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{j}}(r)=\delta_{ij}\int\!\!d^{4}r\;\phi_{\omega}^{\dagger}(r)\phi_{\omega}(r)=\delta_{ij}\!\!\int d^{4}r|\widetilde{\Phi}(r^{\mu})|^{2},$ (51)
also eq.(47) for the dummy integral, $(\int_{\omega})=1$ and $Z^{\prime}|H^{\prime}|^{2}=Z|H|^{2}$ and eq.(40).
The interaction becomes:
$S_{3}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H^{\prime}|^{2}\biggl(2g^{\prime}{}_{0}^{2}N_{c}\frac{\delta^{4}(r^{\mu})}{16M_{0}^{2}}\frac{1}{\sqrt{N}}\int_{\omega}\phi_{\omega}^{\dagger}(r)\frac{1}{\sqrt{N}}\int_{\omega^{\prime}}\phi_{\omega^{\prime}}(r)\biggr)$
$=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H^{\prime}|^{2}\biggl(\frac{g^{\prime}{}_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)N\phi^{\dagger}_{\omega}(0)\phi_{\omega}(0)\biggr)=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;|H|^{2}\biggl(\frac{g_{0}^{2}N_{c}}{8M_{0}^{2}}\delta^{4}(r)|\widetilde{\Phi}(r^{\mu})|^{2}\biggr).$ (52)
IV Full Action
The full action in the pointlike limit then becomes that of the standard model BEH boson coupled to top quarks with in the unprimed parameters:
$S=\!\!\int\!\!d^{4}X\;\biggl(|DH(X^{\mu})|^{2}-\mu^{2}|H|^{2}+g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (53)
where $S^{\prime}$ contains the free (unbound) top quark action and interactions through coloron exchange.
IV.1 BEH Potential
Note the theory generates the usual “sombrero potential”:
$\mu^{2}|H|^{2}+\frac{1}{2}\lambda(H^{\dagger}H)^{2},\qquad\text{where,}\;\;\;\mu^{2}<0.$ (54)
We extremalize the sombrero potential to obtain the broken phase, i.e., for the vacuum of the SM, we therefore find:
$H^{\prime}(X){\Phi}(r)=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+\frac{h(X,r^{\mu})}{\sqrt{2}}\\ 0\end{array}\biggr),$ (57)
where $h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu})$ is the physical BEH boson and $\pi^{a}(X)$ are Nambu-Goldstone bosons, which are phase factors of $H(X^{\mu})$. The electroweak symmetry is broken spontaneously and the gauge fields absorb the Nambu-Goldstone phase factors and acquire mass in the usual way. The resulting broken phase of the BEH field, $h(X)$ is then the standard model and the canonical normalization of $h(X)$ follows from the $Z$ normalization of the internal field $\widetilde{\Phi}(r)$ in eq.(27).
The Nambu-Goldstone modes have only $X$ dependence, and the neutral component of $H^{\prime}$ is then
$H^{\prime}(X)\Phi=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow v_{weak}\widetilde{\Phi}(r)+h(X)\widetilde{\Phi}(r)/\sqrt{2},$ (58)
where $h(X)$ is the SM BEH boson. The “two-body” field, $h(X)$, is now associated with apparent relative time through $\widetilde{\Phi}(r)$, however, the key feature is that $\widetilde{\Phi}(r)\sim\widetilde{\Phi}(0)+{\cal{O}}(r_{\mu}r^{\mu})+$... is a Lorentz invariant function of $r_{\mu}r^{\mu}$ and makes no reference to a particular $\omega^{\mu}$. This expansion can, in principle, yield a form factor in BEH interactions and/or higher dimension operators as described above. Hence the SM BEH field, $h(X)$ is a collective object and its precise two-body nature is actually blended with the collective vacuum $\Phi$. Likewise, the Nambu-Goldstone bosons are “eaten” by the gauge fields in the usual way, and $\Phi(r^{\mu})$ is integrated out and does not affect the usual combinatorial.
The presence of the collective wave-function $\Phi(r^{\mu})$ is not detectable in the kinetic terms, nor in the interaction in the pointlike limit. We obtain the BEH kinetic term eq.III.1) of the SM :
$+\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}\rightarrow\frac{1}{2}ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\;(\partial h(X))^{2}|\widetilde{\Phi}(r)|^{2}=\frac{1}{2}\int\!\!d^{4}X(\partial h)^{2},$ (59)
where we integrate out $\widetilde{\Phi}$ in eqs.(III.1,59) using eq.(40). In fact, there is no way to discern the compositeness of the BEH field from the kinetic and mass terms (this requires the Yukawa interaction and quartic terms). Moreover, $\langle H\rangle\rightarrow v_{weak}$ with the covariant derivative, $D_{\mu}$, of eq.(80), leads to:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S_{1}=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|DH(X^{\mu})|^{2}|\widetilde{\Phi}(r)|^{2}\biggr)\rightarrow$
$=ZM_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr)|\widetilde{\Phi}(r)|^{2}=\!\int\!\!d^{4}X\biggl(M_{W}^{2}W^{+}W^{-}+\frac{1}{2}M_{Z}Z^{2}\biggr).$ (60)
Hence we generate the $W^{\pm}$ and $Z^{0}$ mass terms in the usual way, and the Nambu-Goldstone bosons have become their longitudinal gauge components.
The usual effective action for the BEH boson $h(X)$ emerges:
$\int d^{4}X\biggl(\frac{1}{2}\partial_{\mu}h\partial^{\mu}h-\frac{1}{2}m_{h}^{2}h^{2}-\sqrt{\frac{\lambda}{2}}|\mu|h^{3}-\frac{\lambda}{8}h^{4}\biggr)+\text{electroweak couplings},$ (61)
where $m_{h}=\sqrt{2}|\mu|=125$ GeV.
The structure of the collective field, $\Phi(r^{\mu})$, can in principle be probed through the extended interactions, but the effects will be suppressed. The Yukawa coupling term in natural top condensation takes the form chris1 ; chris2 in the broken phase:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=\sqrt{2N_{c}J}g^{2}_{0}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r^{\mu})\;\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr)\widetilde{\Phi}(r^{\mu}){+h.c.},$ (62)
where $\psi_{L}=(t,b)_{L}$ and $\psi_{R}=t_{R}$ (and $J=16$ is the Jacobian in passing from coordinates $(x,y)$ to $(X,r)$ chris1 ).
To a good approximation we can take the pointlike limit of the potential, and we have $D_{F}(2r^{\mu})\rightarrow(JM_{0}^{2})^{-1}\delta^{4}(r)$. Hence
$g_{Y}=\hat{g}_{Y}|\widetilde{\Phi}(0)|,\qquad\makebox{where,}\qquad\hat{g}_{Y}\equiv g_{0}^{2}\sqrt{2N_{c}/J}=g_{0}^{2}\sqrt{3/8}.$ (63)
This then leads to the conventional mass term for the top quark and BEH coupling:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=g_{Y}\int\!\!d^{4}X\;[\overline{\psi}_{L}(X)\psi_{R}(X]_{f}\biggl(v_{weak}+\frac{h(X)}{\sqrt{2}}\biggr){+h.c.}.$ (64)
IV.2 Integrating Out $r^{\mu}$ and Higher Dimension Operators
If we go beyond the pointlike limit in the BEH-Yukawa interaction we will have Lorentz invariant $O(1/M_{0}^{2})$ corrections. Before taking the pointlike limit of the Yukawa interaction, as in eq.83), we can rewrite this in the broken phase as:
$\!\!\!\!\!\!\!\!\!\!S_{Y}^{\prime}=m_{top}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{L}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}(M_{0}^{2}D_{F}(2r^{\mu}))\;F(r^{\mu})+h.c.,\qquad F=\frac{\widetilde{\Phi}(r)}{\widetilde{\Phi}(0)}\sim 1+a_{1}r^{2}+$... (65)
Note that,
$\!\!\!\!\!\!\!\!\!\!(D_{F}(2r^{\mu}))(r^{p})\sim\int\!\!\frac{d^{4}q}{(2\pi)^{4}}\frac{((\partial_{q})^{p}e^{iqr})}{q^{2}-M_{0}^{2}}\sim\frac{1}{M_{0}^{2+p}}\delta^{4}(r),$ (66)
implying that the expansion in $r$ corresponds to an expansion in $M_{0}^{-1}$, and derivatives $\partial_{X}^{p}$. The expansion in $r^{\mu}$ in all terms in the integrand generates a tower of assorted Lorentz invariant higher dimension operators such as:
$\frac{m_{t}}{M_{0}^{2}}\biggl(1+\frac{h(X)}{\sqrt{2}v}\biggr)\biggl([\overline{b}_{L}(X)D^{2}t_{R}(X)]+[\overline{t}_{L}(X)D^{2}t_{R}(X)]+...\biggr)+h.c.,$ (67)
where $D^{2}$ is the covariant derivative, including the gluon, $\gamma$, $W^{\pm}$, $Z$ couplings. These operators represent new contact terms and processes such as:
$\overline{t}\rightarrow b+W+(g,\gamma,Z))\qquad\text{and,}\qquad\overline{t}\rightarrow t+(g,\gamma,Z)).$ (68)
These can be probed in decays or in production in, e.g., a lepton collider via:
$(\ell^{+}\ell^{-})\rightarrow(\gamma^{*},Z^{*})\rightarrow t+b+(W,g,\gamma,Z)).$ (69)
Determining the full set of effective operators is straightforward, but beyond the scope of the present paper.
V Summary
We have proposed a nontrivial vacuum for the natural top condensation theory. The vacuum is manifestly Lorentz invariant, composed of a collective Lorentz invariant sum over internal wave-functions, $\Phi=(1/\sqrt{N})\sum_{\omega}\phi_{\omega}(r^{\mu})$. Though each internal wave-function, $\phi_{\omega}$, is independent of relative time, $\tau$, they each have dependence upon the arrow of relative time, $\omega^{\mu}$, and are solutions to the Schrödinger-Klein-Gordon equation with eigenvalue $|\mu^{2}|$. The sum over $\phi_{\omega}$ makes the collective field, $\Phi(r^{\mu})$, Lorenz invariant and completely independent of $\omega^{\mu}$.
The bilocal BEH field in the vacuum is then a minimum of the sombrero potential and takes the form:
$H^{\prime}(X)\Phi(r^{\mu})=H(X^{\mu})\widetilde{\Phi}(r)\rightarrow\exp(i\pi^{a}(X)\tau^{a}/2v_{weak})\biggl(\begin{array}[]{c}v_{weak}\widetilde{\Phi}(r^{\mu})+h(X,r^{\mu})\\ 0\end{array}\biggr),$ (72)
where $v_{weak}=|\mu|/\sqrt{2}$ and the observed BEH boson is $h(X)$ where:
$h(X,r^{\mu})=h(X)\widetilde{\Phi}(r^{\mu}).$ (73)
Much can be done to explore this theory further, including the bilocal formalism itself. For a more complete theory, we note that the $b_{R}$ quark also participates in topcolor in an anomaly-free scheme, e.g., see Topcolor , and this should be adapted to the present dynamics. An additional $Z^{0}{{}^{\prime}}$ interaction can be introduced that makes the $\overline{b}b$ channel sub-critical, hence non-binding. Light fermion masses are presumably then generated in analogy to extended technicolor models Technicolor . So far, we have relied upon the intuition of the 1990’s topcolor scheme, but a more complete natural top condensation theory, including the light particle masses and mixing angles, could be readily formulated.
The theory is therefore testable, mainly by direct discovery of the octet of colorons at $M_{0}\approx 6$ TeV. The theory points toward an $SU(3)\times SU(3)\times SU(2)\times U(1)_{Y}$ gauge structure emerging at the $\sim 10$ TeV scale with likely additional $U(1)^{\prime}s$. The theory also offers sensitive probes of new contact interactions involving $t$- and $b$-quarks, and though flavor mixing there may be induced rare processes involving the other quarks and leptons Technicolor .
Note that bilocality of the wave-function is an important naturalness constraint. One might be tempted to, e.g., “loop” the Yukawa interaction and argue for a problematic large correction to the BEH mass $\propto-\overline{g}_{Y}^{2}N_{c}M_{0}^{2}$. This would lead to the false conclusion that the effective theory is “unnatural.” The loop actually generates an enhancing correction to the bilocal potential, i.e., $\sim g_{Y}^{2}\delta^{4}(2r^{\mu})/M_{0}^{2}\sim g_{Y}^{2}D_{F}(2r^{\mu})$ for large $M_{0}$ chris2 , that leads to a “critical amplification” of the effective 4-fermion coupling, $\overline{g}^{2}_{0}$, where: $\overline{g}_{0}^{2}=g_{0}^{2}\left(1-{g_{0}^{2}N_{c}}/{8\pi^{2}}\right)^{-1}.$ Hence, while $\overline{g}_{0}^{2}$ is supercritical, the underlying topcolor coupling, $g_{0}^{2}$, is smaller and subcritical.
In conclusion, we emphasize this theory is natural and manifestly Lorentz invariant. The approximate scale symmetry near critical coupling provides the custodial symmetry of the small $|\mu|^{2}$. The low energy physics is controlled by the $\Phi(r^{\mu})\sim\phi(r^{\mu})$ wave-function spreading, rather than the renormalization group of BHL . The hierarchy is protected against additive radiative corrections by the bilocality, i.e., there is no “additive quadratic divergence,” but only additive and enhancing renormalizations of the bilocal binding interaction chris2 . We used the source/Legendre-transform methods of Jackiw et.al., Jackiw , to derive the effective semiclassical theory used here, which leads to critical amplification of the potential coupling $\overline{g}_{0}^{2}$ chris2 . The fine-tuning is at the few $\%$ level, and, indeed, this may be the first and only minimally-fine-tuned theory of the BEH boson that is consistent with experiment and testable in the not-too distant future.
Appendix A Orthogonality of $\phi_{\omega}(r)$
Recall that the $\phi_{\omega_{i}}(r^{\mu})$ solutions of the Lorentz invariant integro-differential equation are normalized as in eq.(14):
$1=ZM_{0}^{4}\!\!\int\!d^{4}r\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r)=1=M_{0}^{3}\!\!\int\!(ZM_{0})\omega_{\mu i}dr^{\mu}\;d^{3}r_{\perp}\;\phi_{\omega_{i}}^{\dagger}(r)\phi_{\omega_{i}}(r).$ (74)
The latter Lorentz invariant expression can be evaluated in the rest frame, where $\omega_{\mu}=(1,0,0,0)$:
$1=M_{0}^{3}\!\!\int\!d^{3}r|\phi_{\omega}(\vec{r})|^{2},\qquad\text{where,}\qquad 1=ZM_{0}\int dr^{0}\equiv ZM_{0}T.$ (75)
In the small $|\mu|<\!\!M_{0}$ limit the normalization integral eq.(14) is dominated by large $r$, and we have the large distance solution in the rest frame, $\phi(r)\sim Ne^{-|\mu|r}/r$, hence:
$1=M_{0}^{3}\!\!\int\!4\pi r^{2}dr\frac{N^{2}e^{-2|\mu|r}}{r^{2}}\sim 2\pi\frac{N^{2}M_{0}^{3}}{\mu},\qquad\text{hence,}\!\!\qquad\!\!N^{2}=\mu/2\pi M_{0}^{3}.$ (76)
If we consider $Z\sim 1/M_{0}T<\!\!<1$, then $M_{0}\int\omega_{\mu}dr^{\mu}\sim M_{0}T>\!\!>1$, then the $\phi_{\omega}^{\dagger}(r^{\mu})$ become orthogonal in $\omega$,
$\int d^{4}r\;\phi^{\dagger}_{\omega}(r)\phi_{\omega^{\prime}}(r)=0,\qquad\omega_{\mu}\neq\omega^{\prime}_{\mu}.$ (77)
To see orthogonality, consider the timelike hyperboloid defined by $\omega^{2}=1$ and choose $\omega_{0}=(1,0,0,0)$, and $\omega^{\prime}=(\cosh\theta,\sinh\theta,0,0)$ where $\theta$ defines a boost in the $x$ direction. Then $\phi_{\omega_{0}}(r^{\mu})\rightarrow\phi(0,r_{x},r_{y},r_{z})$, with $r^{0}$ as the flat direction for $\omega_{0}$, and $\vec{r}=(r_{x},r_{y},r_{z})$. Then,
$\phi_{\omega^{\prime}}(r^{\mu})=\phi(\omega^{\prime\mu}\omega^{\prime}_{\nu}r^{\nu}-r^{\mu})=\phi(r^{0}\sinh^{2}\theta,r_{x}\cosh^{2}\theta,r_{y},r_{z}).$ (78)
We consider small $\theta$, hence $\phi_{\omega^{\prime}}(r^{\mu})\approx\phi(r^{0}\theta^{2},\vec{r})$. The overlap integral is dominated by the large $r=|\vec{r}|$ component and in this limit with eq.(75), and flat direction $r^{0}$,
$1=ZM_{0}^{4}\!\!\int\!d^{4}r\;\phi_{\omega_{0}}^{\dagger}(r)\phi_{\omega^{\prime}}(r)=ZM_{0}^{4}\!\!\int\!2\pi r^{2}drdr^{0}\frac{N^{2}e^{-2|\mu|r}}{\sqrt{{r}^{2}((r^{0})^{2}\sinh^{2}\theta+{r}^{2})}}$
$\approx\frac{\pi Z{N^{2}}M^{4}_{0}}{\theta\mu^{2}}\ln(M_{0}T)=\frac{1}{2\mu\theta T}\ln(M_{0}T).$ (79)
In the $T\rightarrow\infty$ limit this approaches zero.
The result is not identically zero. In the vacuum, however, where these fields will be clustered into a stable collective state and the system cannot decay, then $T$ can go to infinity with impunity.
Appendix B Brief Summary of Natural Top Condensation
Our formalism, “Natural Top Condensation,” chris1 ; chris2 , is Lorentz invariant and postulates an attractive “topcolor” interaction Topcolor of strength $g_{0}^{2}$ at a high scale $M_{0}$. The bound states are correlated pairs $\overline{\psi}(y)_{L}\psi_{R}(x)\rightarrow\Phi(x,y)\sim\Phi(X,r)$, and we follow Yukawa in writing invariant kinetic terms for the pairs, consistent with their single particle kinetic terms Yukawa . This yields a bound state given by a Schrödinger-Klein-Gordon (SKG) equation satisfied by an internal wave-function, $\phi(r)$. This has eigenvalue $\mu^{2}$, which is the Lagrangian mass of the BEH boson and we have a Yukawa interaction between the bound state and unbound fermions. For supercritical coupling, $g_{0}>g_{c}$ we find $\mu^{2}<0$, and which implies spontaneous symmetry breaking. For small $|\mu|<M_{0}$, near critical coupling, we have significant wave-function spreading and “dilution” of $\phi(0)\sim\sqrt{|\mu|/M_{0}}$. The resulting top quark Yukawa, $g_{Y}\propto\phi(0)$, and quartic couplings, $\lambda\propto|\phi(0)|^{4}$, are subject to power law suppression, rather than the relatively slow renormalization group (RG) evolution in the old Nambu–Jona-Lasinio (NJL NJL ) based top condensation model. The dilution effect significantly reduces the hierarchy and, remarkably, the standard model (SM) quartic coupling, $\lambda\approx 0.25$, becomes concordant with experiment. The fine tuning of the model is also vastly reduced by dilution to $\sim\phi(0)^{2}\sim|\mu|/M_{0}\sim$ few $\%$. Our central prediction is the existence of a binding interaction due to a color octet of massive gluon-like objects, called “colorons,” Topcolor NSD colorons , with mass $M_{0}\sim 6$ TeV. The colorons may be accessible to the LHC. Moreover, loop effects enhance the binding in the $0^{+}$ channel, and the requisite $\mu^{2}<0$ can occur for significantly weaker coloron coupling chris2 . The top condensation model of a composite BEH boson therefore becomes a compelling theory. Inputting the induced Yukawa coupling $g_{Y}\approx 1$ we obtain the resulting prediction $M_{0}\sim 6$ TeV. This construction was confirmed by applying the formal source/Legendre-transformation methods of Jackiw, et. al, chris2 Jackiw .
Starting with third generation fermions, $\psi_{L,R}$, coupled to a coloron exchange potential, we obtain an effective, Lorentz invariant interaction structure for the bilocal BEH boson $H(X^{\mu})\phi(r^{\mu})$ in the symmetric phase of the standard model (SM) chris1 . This was independently derived using techniques of Jackiw, and Cornwall, Jackiw and Tomboulis, chris2 Jackiw . The kinetic terms follow by electroweak gauge invariance, and in ref.chris1 we introduced Wilson lines to “pull-back” the electroweak gauging to $X$. Hence the covariant derivative of $H$ becomes the standard BEH form:
$D_{H\mu}=\frac{\partial}{\partial X^{\mu}}-ig_{2}W^{A}(X)_{\mu}\frac{\tau^{A}}{2}-ig_{1}B(X)_{\mu}\frac{Y_{H}}{2}.$ (80)
With the pullback, $\phi(r)$ is a dimensionless complex scalar and has no gauge charges. The Wilson line pullback is essentially a low energy approximation for the electroweak interactions, but makes the effects of symmetry breaking transparent. - In the barycentric coordinates we have:
$\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!S=M_{0}^{4}\!\!\int\!\!d^{4}X\;d^{4}r\biggl(Z|D_{H}H(X^{\mu})|^{2}|\phi(r)|^{2}+Z|H(X^{\mu})|^{2}|\partial_{r}\phi(r)|^{2}+2{g_{0}^{2}N_{c}}D_{F}(2r)|H^{\dagger}H||\phi(r))|^{2}\biggr)\!\!+S_{Y}+S_{\lambda}+$... (81)
where the $H(X^{\mu})$ kinetic term is canonical with the Lorentz invariant normalization of $\phi(r)$ is:
$1=M_{0}^{4}Z\int\!\!d^{4}r|\phi(r^{\mu})|^{2}\rightarrow M_{0}^{3}\int\!\!d^{3}r|\phi(r^{\mu})|^{2},$ (82)
where $Z$ is as defined upon removal of relative time as in (12). The Yukawa interaction is also generated at tree level,
$\!\!\!\!\!\!\!\!\!\!S_{Y}=\widehat{g}_{Y}M_{0}^{2}\int\!\!d^{4}Xd^{4}r\;[\overline{\psi}_{iL}(X\!+\!r)\psi_{R}(X\!-\!r)]_{f}D_{F}(2r)\;H^{i}(X)\phi(\vec{r}){+h.c.},$ (83)
and a quartic interaction is generated at loop level, given in the point-like $\phi(r)$ approximation by:
$\!\!\!\!\!\!\!\!\!\!S_{\lambda}=-\frac{\widehat{\lambda}}{2}\int\!\!d^{4}X(H^{\dagger}H)^{2}|\phi(0)|^{4}{+h.c.}.=-\frac{\lambda}{2}\int\!\!d^{4}X(H^{\dagger}H)^{2}{+h.c.}.$ (84)
In the above, the Yukawa and quartic couplings $g_{Y}$ and $\lambda$ are derived quantities from the underlying theory. In a pointlike approximation for the interactions, $D(2r)\sim\delta^{4}(r)/M_{0}^{2}$, we have.
$g_{Y}\approx g_{0}^{2}\sqrt{2N_{c}/J}\;\phi(0),\qquad\qquad\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda)\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right).$ (85)
Note that $g_{Y}$ is classical and $\lambda$ arises at loop level (${\cal{O}}(\hbar)$). We thus see that these are subject to dilution in an extended solution with the internal wave-function $\phi(r)$ and $g_{Y}\propto\phi(0)$ and $\lambda\propto|\phi(0)|^{4}$. Experimentally we have $g_{Y}\approx 1$ and $\lambda\approx 0.25$.
The full action then takes the form:
$S=\!\!\int\!\!d^{4}X\;\biggl(|D_{H}H(X^{\mu})|^{2}+|H(X^{\mu})|^{2}S_{\phi}+\ g_{Y}H^{i\dagger}(X)[\overline{\psi}_{R}(X)\psi_{iL}(X)+h.c.]_{f}-\frac{1}{2}\lambda(H^{\dagger}H)^{2}\biggr)+S^{\prime},$ (86)
where [][...] denotes color indices are contracted, and $i$ is an $SU(2)_{weak}$ index. Here $S_{\phi}$ describes the internal wave-function field $\phi(r^{\mu})$, and $S^{\prime}$ describes the coupling of the bound state to external free fermions:
$\!\!\!\!\!\!\!\!\!\!S_{\phi}$ $=$ $M_{0}^{4}\int\!\!d^{4}r\biggl(Z\partial_{\mu}\phi^{\dagger}(r)\partial^{\mu}\phi(r)+2g_{0}^{2}N_{c}D_{F}(2r^{\mu})\phi^{\dagger}(r)\phi(r)\biggr)$ (87)
$\!\!\!\!\!\!\!\!\!\!S^{\prime}$ $=$ $\!\int\!d^{4}x\biggl([\overline{\psi}_{L}(x)iD\!\!\!\!/\!\,\,\psi_{L}(x)]_{f}+[\overline{\psi}_{R}(x)\;iD\!\!\!\!/\!\,\,\psi_{R}(x)]_{f}\biggr)+g_{0}^{2}\!\int\!d^{4}xd^{4}y\;[\overline{\psi}^{i}_{L}(x)\psi_{R}(y)]_{f}D_{F}(x-y)[\overline{\psi}_{R}(y)\psi_{iL}(x)]_{f}.$ (88)
In $S^{\prime}$ we have free unbound fermions, $\psi_{Lif}\sim(t,b)_{L}$ and $\psi_{Rf}\sim t_{R}$ (the minimal model omits $b_{r}$ but this can be readily incorporated as in Topcolor ). Note that the internal field action $S_{\phi}$ is nested within the action for a conventional pointlike BEH boson, $H(X^{\mu})$. $D_{F}(2r^{\mu})$ is the Feynman propagator function for the massive colorons:
$D_{F}(x-y)=-\int\frac{1}{q^{2}-M_{0}^{2}}e^{2q_{\mu}r^{\mu}}\frac{d^{4}q}{(2\pi)^{4}},$ (89)
where $r^{\mu}$ is a radius hence the factor of $2r^{\mu}$. Note the BCS-like enhancement factor of $N_{c}$ in eq.(87)) in the $\phi^{\dagger}D_{F}(2r^{\mu})\phi$ interaction term.
The value of $M_{0}$ is then determined by inputting $g_{Y}=1$, and the known value of the symmetric phase (Lagrangian mass) of the BEH boson, which is $|\mu|^{2}=-(88)^{2}$ GeV2. We find that the scale $M_{0}$ is predicted, $M_{0}\approx 6$ TeV and is no longer the nonsensical $10^{15}$ GeV in the old top condensation based upon the NJL-model BHL . This is due to the faster power-law running of $g_{Y}\propto\phi(0)\sim\sqrt{|\mu|/M_{0}}$, rather than the slow, logarithmic RG running of $g_{Y}$ in the NJL model.
Moreover, a stunning result of this model is the quartic coupling $\lambda$. Experimentally, in the SM using the value of $m_{BEH}\approx 125$ GeV and $v_{weak}\approx 175$ GeV we find $\lambda\approx 0.25$. In the present bilocal scheme, owing to dilution of $\phi(0)$, the quartic coupling is also suppressed and is now generated in RG running from a value of $\lambda(M_{0})=0$ at $M_{0}\approx 6$ TeV, down to $\lambda(|\mu|)$ with $|\mu|\sim 88$ GeV, using $g_{Y}\approx 1$. The prefactor at one loop level reflects the full RG running of $\lambda$, and at leading log the RG equation for $\lambda$ yields RGlam :
$\lambda\approx(g_{Y}^{4}-g_{Y}^{2}\lambda-[\lambda^{2}])\frac{N_{c}}{4\pi^{2}}\ln\left(\frac{M_{0}}{\mu}\right)\approx 0.23\;\;\;\makebox{(cf., $0.25$ experiment.)},$ (90)
where we solve for $\lambda$ self-consistently with $g_{Y}=1$ and $M_{0}\sim 6$ TeV. Note that the $\lambda^{2}]$ term should be omitted since it involves internal propagation in loops of the composite BEH boson (and only slightly affects the result). The $g_{Y}^{2}\lambda$ terms are fermion loop leg renormalizations. This is in excellent agreement with experiment at one loop precision and significantly contrasts the prediction of the old NJL-based top condensation model where the quartic coupling was determined by the RG and found to be $\lambda\sim 1$ BHL , much too large.
The degree of fine-tuning of the theory is remarkably suppressed by $|\phi(0)|^{2}$ in a subtle way. Rather than the naive result one would expect from the NJL model, $\delta g_{0}^{2}/g_{c}^{2}\sim|\mu|^{2}/M_{0}^{2}\sim 10^{-4}$, we now obtain a linear relation: $\delta g_{0}^{2}/g_{c}^{2}\sim|\phi(0)|^{2}\sim|\mu|/M_{0}\sim 1\%$. This is derived in chris1 , but was accidentally noticed when the bound state was treated by a variational “spline approximation” in earlier papers chris2 .
Acknowledgments
I thank Bill Bardeen, Bogdan Dobrescu, Tao Han and other participants of Photon-Lepton 2025, for discussions and comments. Sadly, Bill Bardeen passed away during the completion of this work. His awesome intellect, depth of understanding of physics, and close friendship, are sorely missed.